Assume you put $1,000 into an account today at a bank and leave it there for five years. The bank promises that there will be $1,140.60 in the account at the end of the five years. What is the stated rate of interest if compounded monthly?
Answer #1

We use the formula:
A=P(1+r/12)^12n
where
A=future value
P=present value
r=rate of interest
n=time period.

1,140.60=1000*(1+r/12)^(12*5)

(1,140.60/1000)^(1/60)=(1+r/12)

(1+r/12)=1.00219498

r=(1.00219498-1)*12

=2.63%(Approx).